OpenStudy (anonymous):

What are the x– and y–intercepts of the graph of y = 2x2 – 8x – 10?

OpenStudy (amistre64):

when x=0; the only thing left is y = -10

OpenStudy (amistre64):

when y = 0; its a little trickier, but doable

OpenStudy (amistre64):

0 = 2x^2 -8x -10 ; factor out the 2 ... throw it away, we dont need it 0 = x^2 -4x -5 ; complete the square by +- 4 0 = x^2 -4x +4 -4 -5 ; compress the square and add the constants 0 = (x+2)^2 -9 ; +9 and sqrt the sides +- sqrt(9) = x+2 ; -2 -2 +- 3 = x

OpenStudy (amistre64):

the x ints are: (-5,0) and (-1,0)